OpenStudy (anonymous):

how do you do this ? x+1/3 = 5 - (x+2)/7

OpenStudy (anonymous):

x+1/3=5-(x+2)/7 x- 1/3 - 5 = -(x+2)/7 x + 14/3 = -(x+2)/7 7(x+14/3) = -x-2 7x + 98/3 = -x-2 8x = -2 - 98/3 8x = - 6/3 - 98/3 x = -104 / 64